OpenStudy (anonymous):

the mass of a star is somewhere double without any change in its radius.How does the gravitional field at its surface change?

OpenStudy (anonymous):

F=GM/r^2 F increases proportionally as M increases.

OpenStudy (anonymous):

I think F = GMm/r^2 small m will the mass of the star and big M the mass of the earth.

OpenStudy (anonymous):

You only need to use nayeaddo's equation when you 're calculating the force of gravity BETWEEN two objects. For surface gravity you only need g = GM/r^2. for a star of M = 1, r = 1, g = G. If we double M to get M = 2, g = 2G i.e. the gravitational force at its surface is doubled.

OpenStudy (anonymous):

oh okay I get it